# A Conditional Generative Framework for Synthetic Data Augmentation in Segmenting Thin and Elongated Structures in Biological Images

### v1

@@ 1 Introduction
To address the challenges of acquiring annotated filament data, we propose a conditional generative framework to generate synthetic realistic filaments in microscopy images with paired binary mask as ground truth. Our model is built upon a conditional generative adversarial network[isola2017image], and it learns the mapping between binary filament masks and the corresponding microscopy images. To further enhance the model’s ability to learn the unique structure of filaments, we introduce a filament-aware structural loss that enforces consistency along the filament skeletons. This improves model’s ability to generate realistic connectivity and morphology of filaments. During inference time, the model will take binary masks of curves and generate realistic microscopic images of filaments. This will effectively expand the segmentation dataset and provide a cost-efficient alternative to manual annotation. Different from traditional augmentation methods such as rotation, flipping, perturbations, our generative framework learns the underlying distribution of microscopy appearance. It reproduces similar illumination, texture variations, and noise patterns similar to real fluorescence images while preserving filament geometry and continuity.
@@ 2.1 Filament Segmentation
Segmenting thin and elongated structures is fundamental in biomedical image analysis. Traditional image processing techniques have been widely used for segmenting filamentous structures in biological images [chang, paul, fuller2005filament, almi2015modified, soax, xu20143d, smith2010segmentation]. However, these approaches are highly sensitive to image quality, background noise, and filament density. They often require careful parameter tuning for each dataset. Recent advances in deep learning have substantially improved segmentation performance by learning hierarchical features from data. Deep learning–based segmentation models such as U-Net [ronneberger2015u] and its variants [liu2018densely, ye2025vsr, horiuchi2025deep, liu2019intersection, liu2022extraction] have been successfully applied to segment vessel-like and filamentous structures, achieving better results. Despite the strong performance, deep learning based models heavily rely on large, high-quality annotated datasets. However, creating segmentation datasets for filamentous structures is extremely labor-intensive and costly, as each image can contain hundreds of filaments that need to be traced manually, and may take hours to annotate one image [liu2023pick, liu2018densely]. Moreover, due to the variations in imaging conditions and setups, segmentation models trained on limited dataset often failed to generalize across different imaging settings. Therefore, it is crucial to address the challenge of annotating filaments and enlarge the dataset, generalizing filament segmentation models to diverse imaging settings and biological contexts.
@@ 2.2 Data Augmentation
Imaging processing techniques such as rotation, flipping, and intensity adjustments have been widely used to augment data and improve segmentation performance [chlap2021review, liu2018densely, liu2020quantifying], but these methods only modify existing samples and fail to generate new filament shapes, complex novel patterns, or realistic background variations.
Recent advances in deep generative models allow us to generate realistic synthetic images from learned data distributions. GAN-based [goodfellow2020generative] augmentation has been widely explored to augment limited training data in biomedical image segmentation  [adjei2022examining, motamed2021data]. In [thambawita2022singan], Thambawita et al.proposed SinGAN-Seg to augment polyp images from a single training image. More recently, diffusion models have been a powerful alternative for generative augmentation. Nazir et al. [nazir2025diffusion] introduced a diffusion-based augmentation framework for medical image segmentation that synthesizes anatomically consistent images conditioned on semantic masks. Similarly, the Polyp-DDPM model [dorjsembe2024polyp] proposed diffusion-based model to generate synthetic polyp textures and illumination variations for segmentation tasks.
@@ 3.1 Conditional Generative Model
We employed a conditional generative adversarial network (cGAN) [isola2017image] to model the translation between filament masks and microscopy images. The generator $G$ is used to produce realistic images $y_{\text{gen}}$ conditioned on input masks $x$, while the discriminator $D$ attempts to distinguish between real and generated image–mask pairs. The adversarial objective is defined as:
@@ 3.3 Synthetic Data Generation and Segmentation Training
After training the cGAN on a small set of paired binary masks and real images, we first implemented a random filament mask generator to produce synthetic binary masks. Then we use trained cGAN to generate realistic microscopic images with synthetic binary mask, and creating augmentation dataset. For segmentation network setup, we employ a modified U-Net–based architecture [liu2020quantifying] as the baseline model for its effectiveness in filament segmentation.
@@ 4.1 Datasets
We evaluate the proposed framework on the cytoskeletal filament datasets in [liu2018densely] which contains microtubules and actin filaments. The dataset consists of fluorescence microscopy images captured under different imaging conditions and resolutions with 53 annotated microtubule images of size $1400\times 1524$ and 10 annotated actin filament images of size $1400\times 1524$. We follow the same setup in [liu2018densely], use 25 microtubule images and 2 actin filament images in the datasets for training both cGAN and segmentation model, and testing on the rest 28 microtubule images and 8 actin images.
@@ 4.3 Segmentation Results on Filament Dataset
We adopt the Intersection-over-Union (IoU) and Skeleton IoU (SKIoU) metrics, following the definitions used in [liu2018densely]. IoU measures pixel-level overlap between predicted and ground-truth masks, while SKIoU focuses on structural similarity along filament skeletons.

### v2

@@ 1 Introduction
To address the challenges of acquiring annotated filament data, we propose a conditional generative framework to generate synthetic realistic filaments in microscopy images with paired binary mask as ground truth. Our model is built upon a conditional generative adversarial network[9], and it learns the mapping between binary filament masks and the corresponding microscopy images. To further enhance the model’s ability to learn the unique structure of filaments, we introduce a filament-aware structural loss that enforces consistency along the filament skeletons. This improves model’s ability to generate realistic connectivity and morphology of filaments. During inference time, the model will take binary masks of curves and generate realistic microscopic images of filaments. This will effectively expand the segmentation dataset and provide a cost-efficient alternative to manual annotation. Different from traditional augmentation methods such as rotation, flipping, perturbations, our generative framework learns the underlying distribution of microscopy appearance. It reproduces similar illumination, texture variations, and noise patterns similar to real fluorescence images while preserving filament geometry and continuity.
@@ 2.1 Filament Segmentation
Segmenting thin and elongated structures is fundamental in biomedical image analysis. Traditional image processing techniques have been widely used for segmenting filamentous structures in biological images [3, 17, 6, 2, 22, 21, 19]. However, these approaches are highly sensitive to image quality, background noise, and filament density. They often require careful parameter tuning for each dataset. Recent advances in deep learning have substantially improved segmentation performance by learning hierarchical features from data. Deep learning–based segmentation models such as U-Net [18] and its variants [14, 23, 8, 11, 10] have been successfully applied to segment vessel-like and filamentous structures, achieving better results. Despite the strong performance, deep learning based models heavily rely on large, high-quality annotated datasets. However, creating segmentation datasets for filamentous structures is extremely labor-intensive and costly, as each image can contain hundreds of filaments that need to be traced manually, and may take hours to annotate one image [13, 14]. Moreover, due to the variations in imaging conditions and setups, segmentation models trained on limited dataset often failed to generalize across different imaging settings. Therefore, it is crucial to address the challenge of annotating filaments and enlarge the dataset, generalizing filament segmentation models to diverse imaging settings and biological contexts.
@@ 2.2 Data Augmentation
Imaging processing techniques such as rotation, flipping, and intensity adjustments have been widely used to augment data and improve segmentation performance [4, 14, 12], but these methods only modify existing samples and fail to generate new filament shapes, complex novel patterns, or realistic background variations.
Recent advances in deep generative models allow us to generate realistic synthetic images from learned data distributions. GAN-based [7] augmentation has been widely explored to augment limited training data in biomedical image segmentation  [1, 15]. In [20], Thambawita et al.proposed SinGAN-Seg to augment polyp images from a single training image. More recently, diffusion models have been a powerful alternative for generative augmentation. Nazir et al. [16] introduced a diffusion-based augmentation framework for medical image segmentation that synthesizes anatomically consistent images conditioned on semantic masks. Similarly, the Polyp-DDPM model [5] proposed diffusion-based model to generate synthetic polyp textures and illumination variations for segmentation tasks.
@@ 3.1 Conditional Generative Model
We employed a conditional generative adversarial network (cGAN) [9] to model the translation between filament masks and microscopy images. The generator $G$ is used to produce realistic images $y_{\text{gen}}$ conditioned on input masks $x$, while the discriminator $D$ attempts to distinguish between real and generated image–mask pairs. The adversarial objective is defined as:
@@ 3.3 Synthetic Data Generation and Segmentation Training
After training the cGAN on a small set of paired binary masks and real images, we first implemented a random filament mask generator to produce synthetic binary masks. Then we use trained cGAN to generate realistic microscopic images with synthetic binary mask, and creating augmentation dataset. For segmentation network setup, we employ a modified U-Net–based architecture [12] as the baseline model for its effectiveness in filament segmentation.
@@ 4.1 Datasets
We evaluate the proposed framework on the cytoskeletal filament datasets in [14] which contains microtubules and actin filaments. The dataset consists of fluorescence microscopy images captured under different imaging conditions and resolutions with 53 annotated microtubule images of size $1400\times 1524$ and 10 annotated actin filament images of size $1400\times 1524$. We follow the same setup in [14], use 25 microtubule images and 2 actin filament images in the datasets for training both cGAN and segmentation model, and testing on the rest 28 microtubule images and 8 actin images.
@@ 4.3 Segmentation Results on Filament Dataset
We adopt the Intersection-over-Union (IoU) and Skeleton IoU (SKIoU) metrics, following the definitions used in [14]. IoU measures pixel-level overlap between predicted and ground-truth masks, while SKIoU focuses on structural similarity along filament skeletons.